### v1

# ESPnet-SpeechLM: An Open Speech Language Model Toolkit

Abstract
We present ESPnet-SpeechLM, an open toolkit designed to democratize the development of speech language models (SpeechLMs) and voice-driven agentic applications. The toolkit standardizes speech processing tasks by framing them as universal sequential modeling problems, encompassing a cohesive workflow of data preprocessing, pre-training, inference, and task evaluation. With ESPnet-SpeechLM, users can easily define task templates and configure key settings, enabling seamless and streamlined SpeechLM development. The toolkit ensures flexibility, efficiency, and scalability by offering highly configurable modules for every stage of the workflow. To illustrate its capabilities, we provide multiple use cases demonstrating how competitive SpeechLMs can be constructed with ESPnet-SpeechLM, including a 1.7B-parameter model pre-trained on both text and speech tasks, across diverse benchmarks. The toolkit and its recipes are fully transparent and reproducible at: https://github.com/espnet/espnet/tree/speechlm.
ESPnet-SpeechLM: An Open Speech Language Model Toolkit
Jinchuan Tian1 Jiatong Shi1 William Chen1 Siddhant Arora1 Yoshiki Masuyama2Takashi Maekaku3Yihan Wu1,4Junyi Peng1,5 Shikhar Bharadwaj1Yiwen Zhao1Samuele Cornell1Yifan Peng1 Xiang Yue1Chao-Han Huck Yang6Graham Neubig1Shinji Watanabe1 1Carnegie Mellon University 2Mitsubishi Electric Research Laboratories 3LY Corporation 4Renmin University of China 5Brno University of Technology 6NVIDIA Correspondence: jinchuat@andrew.cmu.edu
1 Introduction
Figure 1: The overview of ESPnet-SpeechLM workflow.
Open-Source Level #Released #Tasks #Tokenizer #Tokenizer #Architectures
Codebase Data Train Infer. Eval. Weights Models Types Choices
VoxtLM Maiti et al. (2024) 1 4 2 2 1
UniAudio Yang et al. (2024b) 1 11 5 5 1
Moshi Défossez et al. (2024) 13 N/A 2 2 1
Mini-Omni Xie and Wu (2024) 1 N/A 2 2 1
GLM-4-Voice Zeng et al. (2024) 3 N/A 2 2 1
ESPnet-SpeechLM (this work) 3 15 10 N/A 4
Table 1: Comparison between ESPnet-SpeechLM and other open-sourced SpeechLM codebases. For open-ended SpeechLM dialogue systems, the #Tasks are not well-defined and left N/A. ESPnet-SpeechLM provides multiple interfaces to bridge a massive number of tokenizer choices and the exact number is also left N/A. Details of the supported features in ESPnet-SpeechLM are in Tab.2. Information as of Dec 2024.
The advent of large language models (LLMs) has significantly advanced machine intelligence, particularly in the text domain Achiam et al. (2023); Dubey et al. (2024). As research expands beyond text, LLMs are increasingly applied to multimodal scenarios Yin et al. (2024); Hurst et al. (2024); Fu et al. (2024), such as speech Cui et al. (2024); Peng et al. (2024a) and vision Zhang et al. (2024a), with the aim of achieving higher-level intelligence and enhancing human-computer interactions. Within this context, Speech Language Models (SpeechLMs) have emerged as a powerful paradigm addressing challenges unique to speech processing.
SpeechLMs have demonstrated remarkable progress across a variety of speech tasks, including zero-shot generalization Wang et al. (2023), low-resource modeling Kharitonov et al. (2023), multi-task learning Maiti et al. (2024); Yang et al. (2024b), instruction following Lu et al. (2024), real-time interaction Défossez et al. (2024); Xie and Wu (2024), and emergent abilities Yang et al. (2024a). Similarly to text-based LLMs Kaplan et al. (2020), SpeechLMs benefit from scaling data volume, parameter size, and computational resources Cuervo and Marxer (2024). These advances have fueled a growing interest in SpeechLM research within the speech and language processing community.
However, despite these advances, the development of SpeechLMs remains a complex and resource-intensive endeavor Défossez et al. (2024). Building such models requires significant expertise and effort across diverse tasks, from data preparation to training, inference, and evaluation. To address these challenges and democratize SpeechLM research, we introduce ESPnet-SpeechLM, an open-source toolkit designed to streamline and accelerate SpeechLM development.
ESPnet-SpeechLM unifies speech tasks under a sequential modeling framework and organizes the SpeechLM development process into a standardized workflow. As illustrated in Fig.1, users begin by defining a custom task template, followed by configuring key parameters. The toolkit then automates all phases of the pipeline: preprocessing, training, inference, and evaluation (§3.2). This modular workflow supports a wide range of design choices, including tokenization methods, model architectures, dynamic multi-tasking, etc. In addition, ESPnet-SpeechLM provides a HuggingFace-compatible interface for sharing datasets and models (§3.3). The toolkit is fully open-source, ensuring reproducibility and accessibility.
To showcase its versatility, we present several use cases demonstrating the scalability and efficiency of ESPnet-SpeechLM. These include building competitive SpeechLM-based automatic speech recognition (ASR) and text-to-speech (TTS) systems on datasets exceeding 200k hours of speech-text paired data (§4.2). We also detail the creation of a 1.7B-parameter multi-task SpeechLM, pre-trained on ASR, TTS, TextLM, and AudioLM tasks, by leveraging 240 billion text tokens or audio frames4.3).
2 Related Work
The ESPnet-SpeechLM toolkit builds upon prior works in two main directions:
Text LLM ecosystem: Some popular development tools in text LLM ecosystems can be generalized to any large-scale sequential modeling task, which means they are also suitable for SpeechLM training. Examples of this include DeepSpeed Rajbhandari et al. (2020) and FlashAttention Dao (2023). To preserve text capability, it is common to initialize SpeechLMs from pre-trained text LLMs, which can rely on open-source platforms like HuggingFace Transformers [^1] https://github.com/huggingface/transformers. These tools are integrated into ESPnet-SpeechLM. We also noticed that current text LLM training frameworks Shoeybi et al. (2019); Zheng et al. (2024) provide limited support for speech features. Our toolkit is presented as a supplement in this direction.
Open-Sourced SpeechLMs and Speech Toolkits: Current research on SpeechLMs and their transparency has been significantly advanced by prior open-source SpeechLM research works Zeng et al. (2024); Xie and Wu (2024); Défossez et al. (2024); Yang et al. (2024b); Maiti et al. (2024). SpeechLM research also greatly benefits from general speech processing and sequence-to-sequence modeling toolkits Watanabe et al. (2018); Ravanelli et al. (2021); Zhang et al. (2024b); Yang et al. (2021); Kuchaiev et al. (2019); Ott et al. (2019), as they provide a wide range of components applicable to SpeechLM development. ESPnet-SpeechLM is presented as a combination of cutting-edge SpeechLM research and well-established speech processing techniques within the open-sourced community. More specifically, it is built upon the existing ESPnet Watanabe et al. (2018) codebase to better exploit prior community efforts and compare with existing non-SpeechLM works. We summarize ESPnet-SpeechLM and related codebases in Tab.1.
3 ESPnet-SpeechLM Toolkit
This section outlines the hierarchical design of the ESPnet-SpeechLM toolkit. We first introduce the fundamental concepts of SpeechLMs in §3.1 followed by a detailed description of the ESPnet-SpeechLM workflow in §3.2. Lastly, we highlight key features of our toolkit in §3.3.
3.1 Speech Language Model
Speech tasks can be generically formulated as predicting target sequences $\mathbf{y}=[\mathbf{y}_{1},...,\mathbf{y}_{N}]$ given input conditions $\mathbf{x}=[\mathbf{x}_{1},...,\mathbf{x}_{M}]$, where each $\mathbf{x}_{m}$ and $\mathbf{y}_{n}$ represents individual data items. $M$ and $N$ stand for the number of data items in conditions and targets, respectively. E.g., for ASR, $\mathbf{x}_{1}$ is the input speech; $\mathbf{y}_{1}$ is the corresponding transcription. Commonly, the training objective is to maximize the posterior $P(\mathbf{y}|\mathbf{x})$.
ESPnet-SpeechLM uniformly frames speech tasks as sequential modeling problems using auto-regressive prediction over discrete token sequences within a decoder-only Transformer Vaswani et al. (2017). Specifically, all data items $\mathbf{x}_{m}$ and $\mathbf{y}_{n}$ are first tokenized into discrete token sequences $\mathbf{x}_{m}^{\text{d}}$ and $\mathbf{y}_{n}^{\text{d}}$. Then, the spliced sequence $\mathbf{s}^{\text{d}}=[\mathbf{x}_{1}^{\text{d}},...,\mathbf{x}_{M}^{\text{d}},\mathbf{y}_{1}^{\text{d}},...,\mathbf{y}_{N}^{\text{d}}]$ serves as the input for model training. Cross-entropy loss optimization over $\mathbf{y}_{1}^{\text{d}},...,\mathbf{y}_{N}^{\text{d}}$ approximates the objective $P(\mathbf{y}|\mathbf{x})$. Predicting $\hat{\mathbf{y}}_{1}^{\text{d}},...,\hat{\mathbf{y}}_{N}^{\text{d}}$ based on the conditions $\mathbf{x}_{1}^{\text{d}},...,\mathbf{x}_{M}^{\text{d}}$ and then detokenizing them into $\hat{\mathbf{y}}_{1},...,\hat{\mathbf{y}}_{N}$ yield the final system prediction.
ESPnet-SpeechLM specifically supports multi-stream language models, i.e., $\mathbf{s}^{\text{d}}\in\mathbb{N}^{T\times n_{q}}$ where $T$ stands for the sequence length and $n_{q}$ stands for the number of streams (See Fig.2). This capability is especially critical for audio codec models Défossez et al. (2022); Zeghidour et al. (2021), which encode each audio frame into multiple tokens. Padding tokens are added to align non-audio data when splicing $\mathbf{s}^{\text{d}}=[\mathbf{x}_{1}^{\text{d}},...,\mathbf{x}_{M}^{\text{d}},\mathbf{y}_{1}^{\text{d}},...,\mathbf{y}_{N}^{\text{d}}]$. These multi-stream models require specialized design considerations discussed in §3.3.
3.2 ESPnet-SpeechLM Workflow
The ESPnet-SpeechLM workflow begins with single-task scenarios and extends naturally to multitask training. In the following, we introduce the concept of the task template (§3.2.1) and describe the end-to-end pipeline from preprocessing to evaluation (§3.2.2-3.2.5). Multitasking is described in 3.2.6.
3.2.1 Task Template
As in §3.1, regardless of the exact sequence $\mathbf{s}^{\text{d}}$, the SpeechLM performs sequential modeling indiscriminately. It is the definition of conditions $\mathbf{x}$, targets $\mathbf{y}$, and the corresponding tokenization methods that give the distinctive composition of $\mathbf{s}^{\text{d}}$ and thus the support of different tasks within SpeechLMs.
Figure 2: The training sequence $\mathbf{s}^{\text{d}}$ is assembled based on the task template, e.g. single-task ASR as depicted here. The sequence is multi-stream with an extra $n_{q}$-axis because the codec tokenizes each frame into multiple tokens.
To handle different speech tasks uniformly, ESPnet-SpeechLM defines each task using a task template, which specifies the composition of the training sequence $\mathbf{s}^{\text{d}}$. As shown in Fig.2, the task template defines the name of the task, the conditions, and the targets. For each data item $\mathbf{x}_{m}$ or $\mathbf{y}_{n}$, the item_name and tokenizer are specified. The training sequence starts from the task identifier, followed by the tokenized sequences from all the data items. For each data item, the first token is a tokenizer indicator; the raw data is tokenized by the specified tokenizer. For single-stream data items and special tokens, padding tokens are added. With this template, the training sequences can be assembled automatically from the given raw data.
3.2.2 Preprocessing
Preprocessing primarily involves tokenization. As some tokenizers are neural network-based and heavy, it is more efficient to conduct the tokenization offline. The tokenization is handled automatically by ESPnet-SpeechLM after receiving a folder for each train/valid/test set as follows. Specifically, an index file is provided for each data item, with the format of example-id content in each line. The name of these index files should correspond to the item_name in the task template.
(folder) train
(file) wav
(line) example-id1 path-to-wav1
(line) example-id2 path-to-wav2
(file) text
(line) example-id1 text1
(line) example-id2 text2
ESPnet-SpeechLM processes these files to generate a unified data.json file for each dataset, which contains tokenized results and metadata. data.json is the data format in both training and evaluation. During preprocessing, all tokenizers in use are detected, and a joint vocabulary is constructed automatically.
Table 2: Summary of supported features in ESPnet-SpeechLM toolkit
Task Templates TextLM, AudioLM, Text-to-Speech, Automatic Speech Recognition, Machine Translation,
Speech-to-Text Translation, Speech-to-Speech Translation, Text-to-Audio, Text-to-Music, Audio Caption,
Text-to-Music, Singing Voice Synthesis, Speech Enhancement, Target Speaker Extraction, Visual TTS
Tokenization Text Subword SentencePiece, HuggingFace Tokenizers
G2P 30 choices
Audio Codec ESPnet-Codec Shi et al. (2024), DAC Kumar et al. (2024),
Encodec Défossez et al. (2022), UniAudio Yang et al. (2024b)
SSL XEUS Chen et al. (2024b), S3PRL Yang et al. (2021),
FairSeq Ott et al. (2019)
Codec_SSL Combine Codec and SSL frame-by-frame
Others Music Score Wu et al. (2024), Vision Token Shi et al. (2022),
Classification Label, Speaker Identity, LLM Embeddings
Modeling & Training Transformer Body ESPnet Builtin, HuggingFace AutoModelForCausalLM
Multi-Stream Vall-E Wang et al. (2023), MultiScale-Transformer Yang et al. (2024b)
Language Model Parallel Interleave, Delay Interleave Copet et al. (2024)
Efficiency DeepSpeed Rajbhandari et al. (2020), FlashAttention Dao (2023), Liger-Kernel
Inference Greedy Search, Beam Search, Top-k Sampling, Top-p Sampling
Evaluation VERSA Shi et al. (2024), with 61 evaluation metrics for speech and audio
Sharing Task Template ESPnet GitHub Repository
Datasets & Models ESPnet HuggingFace Hub
3.2.3 Training
The training behavior of ESPnet-SpeechLM is specified by a configuration file. Besides common training configurations like optimization, batch size, and distributed training setup, the toolkit also supports flexible model architecture configurations for SpeechLM development.
ESPnet-SpeechLM provides multiple implementations of multi-stream language models Wang et al. (2023); Copet et al. (2024); Yang et al. (2024b). The implementations of multi-stream language models all rely on the Transformer body implementation. We provide the ESPnet built-in Transformer implementation to maximize flexibility; alternatively, we support any AutoModelForCausalLM from HuggingFace Transformers to leverage pre-trained text LLMs Also, following Défossez et al. (2024), custom weights can be provided during loss computing to balance the tokens from different tokenization methods. This is usually to guarantee one audio frame has the same loss weight as one non-audio token. Lastly, in addition to applying the cross-entropy loss, the toolkit also supports reinforcement learning from human feedback (RLHF) for SpeechLMs (see Tian et al. (2024b) for details).
3.2.4 Inference
For each of the supported multi-stream language models, we provide multiple inference methods, such as greedy search, beam search, and top-k/top-p sampling. Our implementation also allows multiple heuristics like the min/max generation length. One important heuristic is essential to SpeechLM: unlike text LLMs that only predict text, SpeechLMs need to know the modality of the current predicting target, so that tokens from other modalities can be filtered out to avoid invalid predictions. The current modality is known from the most recent tokenizer indicator (§3.2.1), and will switch when a new tokenizer indicator is predicted.
3.2.5 Evaluation
We create an evaluation script for each supported task. Within these scripts, we consistently adopt the VERSA  [^2] https://github.com/shinjiwlab/versa, a comprehensive collection of >60 speech and audio evaluation metrics Shi et al. (2024). Besides the existing evaluation scripts, a model in a new task can be evaluated simply by specifying the metrics, the inference results, and the reference (if needed).
3.2.6 Multitasking
To build SpeechLMs with versatile functionalities, ESPnet-SpeechLM flexibly supports multitasking. As the SpeechLMs have the same modeling procedure for all tasks, achieving multitasking training is to fuse the training sequences from different tasks in the mini-batches. Similar to single-task training, for each task, the task template definition (§3.2.1) and preprocessing (§3.2.2) are completed separately, which gives multiple tokenized datasets and the corresponding data.json files333 Especially, these preprocessing works are easy to distribute and are suitable for collaborative works. . The data loader accepts a list of data.json and fuses these datasets before training, which allows the users to dynamically change the multitasking data setups. Mini-batches are sampled from the fused datasets during training. Additionally, the sampling ratio among these datasets is adjustable to emphasize some specific data portions.
Table 3: English ASR performance (WER%$\downarrow$) comparison among whisper Radford et al. (2023), OWSM v3.1 Peng et al. (2024b) and ESPnet-SpeechLM ASR (ours). All results are derived from the greedy search.
Whisper OWSM v3.1 ours
small medium small medium
Test sets 244M 769M 367M 1.02B 442M
LS-Clean Panayotov et al. (2015) 3.3 2.8 2.5 2.4 1.9
LS-Other Panayotov et al. (2015) 7.7 6.5 5.8 5.0 4.6
MLS Pratap et al. (2020) 9.1 10.2 8.1 7.1 7.2
TEDLIUM3 Hernandez et al. (2018) 4.6 5.1 5.0 5.1 5.7
WSJ Paul and Baker (1992) 4.3 2.9 3.8 3.5 5.2
FLEURS Conneau et al. (2022) 9.6 6.4 10.3 9.0 7.7
Avg. 6.4 5.7 5.9 5.4 5.4
3.3 Supported Features
We summarize the core configurable features in ESPnet-SpeechLM workflow in Tab.2 and highlight them as follows:
Tokenization: For text, we support subword models and grapheme-to-phoneme (G2P) tools, with an emphasis on HuggingFace tokenizers. For audio tokenization, we support both audio codec models and self-supervised learning (SSL) tokens. We provide multiple options for these two tokenization methods, with an emphasis on ESPnet-Codec Shi et al. (2024) and XEUS Chen et al. (2024b). Additionally, we find that concatenating codec and SSL tokens frame-by-frame behaves well in both speech understanding and generation. Besides text and audio, these multi-modal models can leverage information from auxiliary modalities, such as music score Wu et al. (2024), vision token Shi et al. (2022), classification labels (e.g., bool, time-stamp), speaker-identity and the continuous LLM embeddings.
Training: As in §3.2.3, for the Transformer body, we provide the ESPnet built-in implementation as well as the HuggingFace Transformers implementation. Upon the Transfomer, we support 4 distinctive multi-stream language model implementations Wang et al. (2023); Yang et al. (2024b); Copet et al. (2024). For training efficiency, we leverage DeepSpeed Rajbhandari et al. (2020), FlashAttention Dao (2023) and Liger-Kernel [^4] https://github.com/linkedin/Liger-Kernel. These modules enable us to achieve model FLOPs utility (MFU) Chowdhery et al. (2023) as high as 35% with multi-node training using NVIDIA H100 GPUs.
Inference, Evaluation, and Sharing: For all supported architecture, we provide all 4 inference methods. VERSA provides more than 60 speech-related evaluation metrics. To ensure transparency and reproducibility, the code and task templates are released through the ESPnet GitHub repository; tokenized datasets and pre-trained models are released through ESPnet Huggingface Hub.
Table 4: TTS performance on full LibriSpeech Test-Clean Panayotov et al. (2015). SPK_SIM is measured only when zero-shot speaker prompting is supported. The speaker prompts are the same for all tests. All results from VERSA. No post-selection applied. $\star$ means third-party implementation.
Model WER($\downarrow$) SPK_SIM($\uparrow$) Proxy MOS($\uparrow$)
ChatTTS 2Noise (2024) 7.1 - 3.52
CosyVoice Du et al. (2024) 5.0 0.51 4.15
Parler-TTS Lacombe et al. (2024) 4.7 - 3.83
WhisperSpeech Collabora (2024) 13.5 - 4.06
VallE-X $\star$ Zhang et al. (2023) 27.3 0.35 3.38
VallE 2 $\star$ Chen et al. (2024a) 27.8 0.46 3.65
ESPnet-SpeechLM TTS (ours) 3.1 0.55 4.03
4 User Cases
Table 5: Evaluation on the multitask pre-trained SpeechLM using ESPnet-SpeechLM and its comparison with prior text LLM, SpeechLMs, and Multimodal LMs. The numbers of competitors are from their own report unless marked by $\star$. - means unreported numbers.
Task ASR TTS TextLM AudioLM
Metric Size WER($\downarrow$) WER($\downarrow$) SPK_SIM($\uparrow$) Proxy MOS($\uparrow$) MMLU($\uparrow$) ARC-C($\uparrow$) HS($\uparrow$) OBQA($\uparrow$) Perplexity($\downarrow$)
LLaMA-3.2 Dubey et al. (2024) 1B - - - - 32.2 32.8 41.2 29.2 -
VoxtLM Maiti et al. (2024) 1.3B 2.7 / 6.5 - - - - - - - 40.9
Moshi Défossez et al. (2024) 7B 5.7 / - 0 4.7 - - 49.8 - - - -
MiniOmni Xie and Wu (2024) 0.5B 4.5 / 9.7 - - - - - - - -
VITA Fu et al. (2024) 8x7B 8.1 / 18.4 - - - 71.0 - - - -
GLM-4-Voice Zeng et al. (2024) 9B 2.8 / 7.7 5.6 - - - - - - -
ESPnet-SpeechLM (ours) 1.7B 2.8 / 5.9 6.0 0.701 3.99 30.5 41.3 50.4 31.4 16.4
This section provides several user cases to demonstrate the performance of SpeechLMs built from ESPnet-SpeechLM. We first build single-task SpeechLM-Style ASR and TTS models in §4.2. As a highlight of this demo, we present a 1.7B pre-trained SpeechLM that covers 4 tasks similar to Maiti et al. (2024): ASR, TTS, text auto-regressive prediction (TextLM), and speech auto-regressive prediction (AudioLM) (§4.3). These models are released through ESPnet HuggingFace Hub [^5] https://huggingface.co/espnet.
4.1 Experimental Setups
Model and Tokenization: We consistently leverage the pre-trained text LLM, SmolLM2 series [^6] https://huggingface.co/HuggingFaceTB, for SpeechLM initialization. We adopt the 360M and 1.7B versions for single-task and multi-task models, respectively. We adopt delay interleave Copet et al. (2024) as the multi-stream language model architecture. In terms of tokenization, we adopt the Codec_SSL method (§3.3) for speech representation. To preserve full transparency and self-consistency, ESPnet-Codec [^7] https://huggingface.co/ftshijt/espnet_codec_dac_large_v1.4_360epoch and XEUS [^8] https://huggingface.co/espnet/xeus; K-Means tokenizer trained on its last layer of representation using 5k clusters are adopted for codec and SSL tokenizers respectively.
Data, Training, and Inference: We collect open-sourced data for all experiments. Our data contains 200k hours of speech and 115B tokens of text, most in English. When expanding speech data into ASR, TTS, and AudioLM tasks, this is equivalent to 240B text tokens or audio frames. Detailed data composition is in Appendix A. We balance the weights for text, SSL, and codec tokens as 1: 0.5: 0.0625999Each audio frame is represented by 1 SSL token and 8 codec tokens. This ratio is to ensure (1) the text tokens have the same weight as the audio frames, and (2) SSL tokens have the same weight as 8 codec tokens combined.. The training used 8/24 H100 GPUs for single/multi-task training. We use batch size as large as around 2M frames or tokens and a constant learning rate of 2e-4, with 10k warmup steps. We train the model for 2 data passes. We use greedy search for ASR and top-k sampling for TTS ($k=30,temperature=0.8$).
Evaluation: Following Maiti et al. (2024); Tian et al. (2024b), we test word error rate (WER) for ASR; ASR WER, Speaker Similarty and Proxy MOS for TTS; sWUGGY and sBLIMP Nguyen et al. (2020) for AudioLM. We measure the TextLM ability using popular metrics like MMLU Hendrycks et al. (2020), ARC-Challenge (ARC-C) Clark et al. (2018), HellaSwag (HS)Zellers et al. (2019), and OpenBookQA (OBQA) Mihaylov et al. (2018).
4.2 ASR and TTS Experiments
We evaluate our ASR system on multiple benchmarks and compare it with the popular open-sourced ASR models: whisper-v3-large Radford et al. (2023) and OWSM v3.1-medium Peng et al. (2024b). As suggested in Tab.3, our SpeechLM-based ASR system achieves comparable results in English with these two popular speech recognizers even using much fewer parameters. In Tab.4, we compare the ESPnet-SpeechLM TTS system with other discrete-based TTS systems [^10] For VallE-X and VallE 2, we use the third-party implementations: https://huggingface.co/Plachta/VALL-E-X/resolve/main/vallex-checkpoint.pt, https://huggingface.co/amphion/valle. The results suggest our TTS system achieves decent performance on all evaluation metrics.
4.3 Multi-Task Experiments
We demonstrate the performance of our multitask pre-trained SpeechLM in Tab.5. Compared with other SpeechLMs Maiti et al. (2024); Xie and Wu (2024); Zeng et al. (2024); Fang et al. (2024) and multimodal LMs Fu et al. (2024), our pre-trained model still preserves decent ASR, TTS and AudioLM performance even with limited parameter budget. In terms of text capability, the pre-trained model preserves close performance compared with the text-only LLM LLaMA-3.2-1B Dubey et al. (2024).
5 Future Works
We will continue the development of the ESPnet-SpeechLM toolkit, such as supporting more tokenization methods, more task templates, more modeling options, and LLM inference engines Kwon et al. (2023). We are also interested in applying this toolkit to our SpeechLM research. For pre-training, we are interested in larger-scale models and models that can capture rich paralinguistic information in speech. For post-training, we are interested in achieving conversational interactions, speech-based instruction following ability, and even agent-alike behaviors. Our plan also includes real-time and duplex design, HFRL for SpeechLM and SpeechLMs that trained from flat start.
6 Conclusion
This demo presents ESPnet-SpeechLM, a toolkit that covers the whole workflow of speech language model development, with comprehensive support in multiple design choices. We also provide user cases for both single-task and multi-task training, showing competitive performance with other models in the market. The toolkit promises to keep full transparency in data, code, recipes, and pre-trained models.
Appendix A Data Details
The statistics of our training data are in Tab.6. We highlight as follows.
Speech Data: We collected 213k hours of open-source speech data and applied the following preprocessing. (1) We only use the English subset of Emilia He et al. (2024); (2) We use the Emilia pipeline He et al. (2024) to process the raw audio files in the English subset of Yodas Li et al. (2023); and (3) We only use the English subset of the OWSM Tian et al. (2024a) dataset. We exclude the MLS to avoid duplication. This data is not applied to TTS as the speaker identity is absent.
Text-Only Data: The text pretraining dataset is a diverse and extensive collection of text data sourced from three primary domains, encompassing a total of 115.69 billion tokens. The largest segment, contributing 82.36 billion tokens, is derived from general web content (FineWeb-EDU [^11] https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu ), offering a rich variety of information spanning numerous topics and styles, suitable for broad language understanding tasks. Complementing this is 20.56 billion tokens of multilingual text from the Multilingual CC News dataset [^12] https://huggingface.co/datasets/intfloat/multilingual_cc_news, which enhances the model’s ability to comprehend and generate text across multiple languages, catering to global linguistic diversity. Lastly, 12.77 billion tokens are sourced from the OpenCoder Annealing Corpus Huang et al. (2024), a code-centric dataset, which bolsters the model’s proficiency in understanding and generating programming languages and technical instructions. Together, these datasets provide a balanced blend of general, multilingual, and technical data, creating a robust foundation for versatile language model capabilities.
Table 6: Detailed composition of the training data used in this work
Dataset #Hours Text Tokens or Audio Frames (B)
ASR TTS TextLM AudioLM
LibriSpeech Panayotov et al. (2015) 960 0.18 0.33 - 0.17
MLS Pratap et al. (2020) 55k 9.88 16.13 - 9.15
Emilia He et al. (2024) 50k 9.10 18.16 - 8.33
Yodas Li et al. (2023) 80k 12.14 23.75 - 11.14
OWSM Tian et al. (2024a) 27k 5.29 - - 4.95
General Text 82.36
Multilingual Text 20.56
Code 12.77
Total 213k 36.59 58.37 115.69 33.74